Circular No. 6171

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

GRS 1915+105 K. Ebisawa and N. E. White, Goddard Space Flight Center, NASA;

T. Kotani, Institute of Space and Astronautical Science; and A. Harmon, Marshall Space Flight Center, NASA, communicate: “The Galactic superluminal jet source GRS 1915+105 (Mirabel and Rodriguez 1994, Nature, 371, 46) was observed to be unusually bright with the ASCA satellite on Apr. 20.4–20.9. The flux in the band 2–10 keV was ∼ 5 × 10^{−9} erg cm^{−2} s^{−1} or 0.26 Crab. The spectrum is well fit by a power-law with photon index 1.4 between 1 and 5 keV, and it is significantly absorbed by cold matter corresponding to a hydrogen column density of 5 × 10^{22} cm^{−2}. Above 5 keV, the spectrum decays exponentially with an e-folding energy of 5 keV. The near simultaneous BATSE flux in the band 20–100 keV was < 50 mCrab, confirming the strong high-energy cut-off. The ASCA spectrum and the BATSE flux are similar to those in the 1994 Sept. observations, and the overall spectrum seems much softer than that during the 1992–1993 outburst, when the source was bright in both soft and hard x-rays. Follow-up x-ray and radio observations are encouraged.”

V1974 CYGNI V. Elkin, Special Astrophysical Observatory, writes: “CCD spectro-

grams (range 649–661 nm, resolution 0.075 nm) of V1974 Cyg obtained on 1994 Oct. 15–16 with the 6-m telescope (+ Main Stellar Spectrograph) show a broad Hα emission line split into several components. The central narrow emission feature (FWHM = 0.1 nm) may belong to a vast diffuse cloud larger than 33′′ (which is the length of the spectrograph slit); this feature has radial velocity v_{r} = −19.2 ± 1.0 km/s, which is near that of the interstellar Ca II 396.8-nm line (v_{r} = −25.8 ± 1.4 km/s) as seen in the early nebular stage; such a cloud may be interstellar and thus not connected with the nova eruption. A Gaussian analysis yielded the relative intensities of Hα to N II 654.8-nm to N II 658.4-nm as 100:10:30.”

NOTICE CONCERNING TELEGRAMS Because of the steady decline in telegram communications in recent

years, with very few telegram messages received by the Central Bureau in recent months, it is planned that the Bureau’s TWX number (710-320-6842) will be retired as of 1995 July 1. The Bureau ceased transmitting telegram messages in 1993 (cf. IAUC 5873).

1995 May 11 Daniel W. E. Green